# The Unbearable Slowness of Being

### v1

@@
The Unbearable Slowness of Being
@@ 1 Abstract
This article is about the neural conundrum behind the slowness of human behavior. The information throughput of a human being is about 10 bits/s. In comparison, our sensory systems gather data at an enormous rate, no less than 1 gigabits/s. The stark contrast between these numbers remains unexplained. Resolving this paradox should teach us something fundamental about brain function: What neural substrate sets this low speed limit on the pace of our existence? Why does the brain need billions of neurons to deal with 10 bits/s? Why can we only think about one thing at a time? We consider plausible explanations for the conundrum and propose new research directions to address the paradox between fast neurons and slow behavior.
@@ 2 The information rate of human behavior
“Quick, think of a thing… Now I’ll guess that thing by asking you yes/no questions.” The game ‘Twenty Questions’ has been popular for centuries [1] as a thinking challenge. If the questions are properly designed, each will reveal 1 bit of information about the mystery thing. If the guesser wins routinely, this suggests that the thinker can access about $2^{20}\approx 1$ million possible items in the few seconds allotted. So the speed of thinking – with no constraints imposed – corresponds to 20 bits of information over a few seconds: a rate of 10 bits/s or less.
Using the tools from information theory [2, 3] one can evaluate the bit rate of human actions and thoughts in many different circumstances based on a common currency. The basic approach is to assess the range of possible actions that a person may execute in a given time. Along the way, one needs a clear criterion to distinguish the action from its noisy variations. This distinction of signal and noise is quantified by Shannon’s entropy and ultimately leads to an information rate, expressed in bits/s (see Appendix A).
Take for example a human typist working from a hand-written manuscript. An advanced typist produces 120 words per minute [dhakal_observations_2018]. If each word is taken as 5 characters, this typing speed corresponds to 10 keystrokes a second. How many bits of information does that represent? One is tempted to count the keys on the keyboard and take the logarithm to get the entropy per character, but that is a huge overestimate. Imagine that after reading the start of this paragraph you are asked what will be the next let…
English contains orderly internal structures that make the character stream highly predictable. In fact, the entropy of English is only $\sim$ 1 bit per character [shannon_prediction_1951]. Expert typists rely on all this redundancy: If forced to type a random character sequence, their speed drops precipitously. So we conclude that the typist operates at
@@
One might argue that the typist is hampered by the speed of the fingers. Perhaps the information rate would be greater when using a motor system that evolved for communication, like the human voice. For the audience to comfortably follow an instructional presentation in English, the recommended narration rate is 160 words per minute [williams_guidelines_1998]. That is indeed slightly higher than the 120 words/min typing rate, yielding an information rate of 13 bits/s.
@@
In Appendix B and Table 1 we review additional measurements of human behavioral throughput, ranging from classic laboratory experiments to recent E-Sports competitions. These studies span the better part of a century with a remarkably concordant result: Humans operate at an information rate of about 10 bits/s, and the rates of sensory perception and motor output appear evenly matched.
How should one interpret that number? The most startling aspect is how ridiculously small it is compared to any information rate we encounter in daily life. For example, we get anxious when the speed of our home WiFi network drops below 100 Mbit/s because that might compromise our enjoyment of Netflix shows. Meanwhile, even if we stay awake during the show, our brain will never absorb more than 10 bits/s of that giant bitstream. More relevant to the present arguments, the speed of human behavior is equally dwarfed by the capacity of neural hardware in our brains, as elaborated in the following section.
@@ 3 The information capacity of the nervous system
From the moment that Claude Shannon founded information theory [shannon_mathematical_1948], neurophysiologists have applied that framework to signaling in the nervous system. For example, one can view a single neuron as a communication channel. In sensory systems, in particular, we are quite sure what the neurons are communicating about, namely the sensory stimulus, which also happens to be under the experimenter’s control. To measure how much information the neuron conveys, one proceeds as described above for behavior: (1) Identify the range of outputs of the neuron across all possible stimuli. (2) Determine what part of that range is “noise” vs “signal”; typically this involves repeating the same stimulus many times and measuring the response variation. (3) Use the expressions for entropy and mutual information to derive an information rate (see Appendix C).
For example, the photoreceptor cells in the human eye convert dynamic light input to a continuously varying membrane voltage. Using the approach described in Appendix C, one finds that one human cone photoreceptor can transmit information at $\approx 270$ bits/s. The population of 6 million cones in one eye has a capacity of about 1.6 gigabits/s.
Unlike photoreceptors, most neurons in the central nervous system communicate via action potentials. The message from one neuron to the next depends entirely on the timing of their spikes. Again, one can adapt the methods of information theory and ask how much information gets transmitted that way. The answer is about 2 bits per spike [borst_information_1999]. This holds across neurons from many different species and over a wide range of firing rates.
This ballpark number of 2 bits/spike allows us to estimate the information rate at different stages of sensory processing. For example, the output of the eye consists of 1 million axons from retinal ganglion cells [kandel_low-level_2021]. With particularly strong stimuli one can drive them at an average rate of 50 Hz. So the capacity of the optic nerve is about 100 megabits/s or less. Note this is 10 times smaller than the capacity of cone photoreceptors: The circuits inside the retina already compress the visual signal substantially, by at least a factor of 10.
The example of retinal ganglion cells suggests that when driven to high firing rates, individual neurons may transmit hundreds of bits per second. In the mammalian cortex, the average firing rate of a neuron is very low, typically a few spikes per second [lennie_cost_2003]. However, even under those conditions, a single neuron will transmit $\sim$10 bits/s – equivalent to the information throughput of the entire human being.
@@ 4 The paradox of slow behavior
So why do we need such an enormous number of neurons – many billions of them in the cerebral cortex alone – to implement the slow pace of behavior? To reiterate the problem: Human behaviors, including motor function, perception, and cognition, operate at a speed limit of 10 bit/s. At the same time, single neurons can transmit information at that same rate or faster. Furthermore, some portions of our brain, such as the peripheral sensory regions, clearly process information at least a million-fold faster. Some obvious questions arise: What sets the speed limit on human behavior? And what can that teach us about the neural mechanisms of cognition?
Before attempting an answer, we should agree on the magnitude of the question, which sets an expectation for the magnitude of the answer. We need to explain a factor of a million (conservatively), namely the ratio between peripheral information processing and the overall throughput of the brain. Numerically speaking, this may be the largest unexplained number in brain science. This order of magnitude sets a filter for hypothetical solutions. Suppose, for example, that a particular hypothesis explains why people process information two times faster in one task condition than in another. That might be enough for a paper in a prominent journal, but the tiny effect size doesn’t even make a dent in the paradox. Obviously we should move on and read a different article.
@@ 4.1 What about photographic memory?
One scientific report stands out here, about a single human subject who could fuse a random dot stereogram after each monocular image had been presented on different days [stromeyer_detailed_1970]. Each image had $100\times 100$ pixels and was presented for 2 minutes, which would require a perceptual rate of 100 bits/s, ten times what we propose. Now a ten-fold discrepancy is nothing dramatic, given the million-fold conundrum we pose. Nevertheless, there are good reasons to question this claim. For example, neural signals from the two eyes are merged already in the primary visual cortex. So the subject’s brain would have to somehow inject the memorized image into neurons very early in the visual system, or perhaps even store it there. The claim is so unexpected and potentially revolutionary that it should have triggered many follow-up experiments. No supporting reports have appeared, suggesting that those studies did not go as hoped for.
@@ 4.2 But what about the rich detail everywhere in my visual scene?
However, this is an illusion, called “subjective inflation” in the technical jargon [oregan_solving_1992, knotts_subjective_2019]. People feel that the visual scene is sharp and colorful even far in the periphery because in normal life we can just point our eyes there and see vivid structure. In reality, a few degrees away from the center of gaze our resolution for spatial and color detail drops off drastically, owing in large part to neural circuits of the retina [banks_peripheral_1991]. You can confirm this while reading this paper: Fix your eye on one letter and ask how many letters on each side you can still recognize [rayner_eye_1998]. Another popular test is to have the guests at a dinner party close their eyes, and then ask them to recount the scene they just experienced. These tests indicate that beyond our focused attention, our capacity to perceive and retain visual information is severely limited, to the extent of “inattentional blindness” [simons_gorillas_1999, cohen_what_2016].
@@ 4.3 But what about unconscious processing?
Another objection to a very low rate of perception is that, by definition, it only takes into account things that we can consciously experience and access so that we can talk about them [graham_what_1869]. Surely, this argument goes, the myriad images that flit over our retina every day must carry a great deal of fine-grained information into the brain. Perhaps this information cannot be accessed consciously as photographic memory, but it is still stored in the brain. In its most general formulation, this idea cannot really be falsified: We can only check whether an animal has information by experimenting on its behavior, and that behavioral output is limited to 10 bits/s over the lifetime of the animal. So one cannot directly test whether the brain can acquire information above the rate at which it can be read out.
@@
A specific proposal for unconscious processing involves the development of visual circuits. It is claimed that the visual system learns about the statistics of natural scenes from being exposed to millions of those images over time. Through mechanisms of synaptic plasticity [bi_synaptic_2001], the structure of neural circuits responds to visual experiences in the course of early development, which alters the subsequent function of those circuits [katz_synaptic_1996]. How much information does this process extract from the environment?
@@
The famous stripe-rearing experiments on kittens can serve as an example [stryker_physiological_1978]. The kittens were raised in an environment painted with vertical black and white stripes. From the kitten’s perspective, because it still got to shake its head, the orientation of the stripes was restricted to perhaps $\pm$20 degrees from the vertical. This did lead to some reorganization of brain wiring in the visual cortex. Let us suppose that the kitten brain has completely incorporated the change in visual statistics. Whereas in the natural world, all orientations are about equally represented across 360 degrees, in the new environment they are restricted to a 40-degree range. The information associated with such a change in probability distributions is $\log_{2}(360/40)\approx 3$ bits.
@@
By perfectly learning the new distribution of angles, the kitten acquired only three bits. The same argument applies to other aspects of natural image statistics: The distribution of spatial and temporal frequencies differs from the uniform, but the shape of that distribution can be captured by just a few bits [olshausen_natural_1996]. This is also true of the distribution of color spectra. In each of these cases, processes of activity-dependent development can learn only a few bits over the animal’s lifetime.
@@
Why is the information about scene statistics so limited? The main reason is that they are translation-invariant: To good approximation the frequency spectrum of a natural image is the same in the bottom left as in the top right. Consequently, it is not necessary to learn separate statistics at each point in the visual field or to specify neural circuitry differently at each point. Nature “knows” about this translation-invariance and has exploited it in a clever way: by inventing families of neuronal cell types that are cloned across the visual field [roska_retina_2014, sanes_types_2015]. For example, the human retina contains many millions of neurons, but they come in only $\sim 100$ types [grunert_cell_2020]. Cells of a given type are distributed evenly across the retina. And so are the synaptic circuits between cells, because synapses are specified by type identity. In this way, modifying a single gene can alter synaptic circuitry in the same way at thousands of different sites within the visual field [sanes_synaptic_2020]. Modern image recognition systems have copied this invention in the form of convolutional neural networks, where all units within a layer share the same convolution kernel, thus requiring only a few parameters to be learned [lecun_gradient-based_1998].
@@
Our nervous system uses this convolutional trick even when adapting in real-time to a change in image statistics. For example, electrical synapses in the retina can be modified by the neuromodulator dopamine [bloomfield_diverse_2009]. Thus a change in a single variable, namely the diffuse concentration of dopamine, modifies visual processing everywhere in the visual field. These adaptations occur on the scale of minutes, so again, only a small rate of information flow from the environment is needed for these updates.

### v2

@@
The Unbearable Slowness of Being: Why do we live at 10 bits/s?
@@
In Brief
Zheng and Meister write about the paradoxical slowness of human behavior. While our senses gather data at $10^{9}$ bits/s, our overall information throughput is only $10$ bits/s. This stark contrast touches on many fundamental aspects of brain function.
@@ 1 Abstract
This article is about the neural conundrum behind the slowness of human behavior. The information throughput of a human being is about $10$ bits/s. In comparison, our sensory systems gather data at $\sim 10^{9}$ bits/s. The stark contrast between these numbers remains unexplained and touches on fundamental aspects of brain function: What neural substrate sets this speed limit on the pace of our existence? Why does the brain need billions of neurons to process 10 bits/s? Why can we only think about one thing at a time? The brain seems to operate in two distinct modes: the “outer” brain handles fast high-dimensional sensory and motor signals, whereas the “inner” brain processes the reduced few bits needed to control behavior. Plausible explanations exist for the large neuron numbers in the outer brain, but not for the inner brain, and we propose new research directions to remedy this.
2 Introduction
“Quick, think of a thing… Now I’ll guess that thing by asking you yes/no questions.” The game ‘Twenty Questions’ has been popular for centuries 1 as a thinking challenge. If the questions are properly designed, each will reveal 1 bit of information about the mystery thing. If the guesser wins routinely, this suggests that the thinker can access about $2^{20}\approx 1$ million possible items in the few seconds allotted. So the speed of thinking – with no constraints imposed – corresponds to 20 bits of information over a few seconds: a rate of 10 bits per second or less.
More generally, the information throughput of human behavior is about 10 bits/s. We review measurements spanning the better part of a century that involve all aspects of human cognition: perception, action, or – as in the example above – imagination. The general approach is to assess the range of possible actions that a person may execute in a given time. Along the way, one needs a clear criterion to distinguish the action from its noisy variations. This distinction of signal and noise is quantified by Shannon’s entropy and ultimately leads to an information rate, expressed in bits/s (see Box).
This information-theoretic approach allows us to compare the speed of processing across different mental tasks and processes, between different neural structures in the same brain, across different species, and between brains and machines. Obviously, this is just one framework by which to characterize human experience, but it offers valuable insights through comparative analysis.
In particular, our peripheral nervous system is capable of absorbing information from the environment at much higher rates, on the order of gigabits/s. This defines a paradox: The vast gulf between the tiny information throughput of human behavior, and the huge information inputs on which the behavior is based. This enormous ratio – about 100,000,000 – remains largely unexplained.
We show that this paradox has implications for many areas of brain science, from neurotheory to assistive technology. A key part of the conundrum is: Why can we only do one thing at a time, whereas our peripheral nervous system processes thousands of items in parallel? We survey various candidate explanations, spanning from evolution to neural circuitry, and find that many of these lead to open questions. We conclude by making suggestions for future neuroscience research to address the conundrum posed by the slowness of being.
@@ 3 The information rate of human behavior
Take for example a human typist working from a hand-written manuscript. An advanced typist produces 120 words per minute dhakal_observations_2018. If each word is taken as 5 characters, this typing speed corresponds to 10 keystrokes a second. How many bits of information does that represent? One is tempted to count the keys on the keyboard and take the logarithm to get the entropy per character, but that is a huge overestimate. Imagine that after reading the start of this paragraph you are asked what will be the next let…
English contains orderly internal structures that make the character stream highly predictable. In fact, the entropy of English is only $\sim$ 1 bit per character shannon_prediction_1951. Expert typists rely on all this redundancy: If forced to type a random character sequence, their speed drops precipitously. So we conclude that the typist operates at
@@
One might argue that the typist is hampered by the speed of the fingers. Perhaps the information rate would be greater when using a motor system that evolved for communication, like the human voice. For the audience to comfortably follow an instructional presentation in English, the recommended narration rate is 160 words per minute williams_guidelines_1998. That is indeed slightly higher than the 120 words/min typing rate, yielding an information rate of 13 bits/s.
@@
In Appendix A and Table 1 we review additional measurements of human behavioral throughput, ranging from classic laboratory experiments to recent E-Sports competitions. These studies, spanning the better part of a century and divergent fields, all point to a remarkably concordant result: Humans operate at an information throughput rate of about 10 bits/s, and the rates of sensory perception and motor output appear evenly matched. As discussed above, all these tasks involve some criteria for a correct solution, which allows one to separate signal from noise. The range covers many types of human behavior, from games to paid professions, but obviously leaves out others (dancing, dreaming,…) for which an information throughput would be hard to define, let alone measure.
Behavior/Activity Time scale Information rate (bits/s) References
Binary digit memorization 5 min 4.9  internationalassociationofmemory_5_2024
Blindfolded speedcubing 12.78 sec 11.8  guinnessworldrecordslimited_fastest_2023
Choice-reaction experiment min $\sim$ 5 hick_rate_1952, hyman_stimulus_1953, klemmer_rate_1969
Listening comprehension (English) min - hr $\sim$13 williams_guidelines_1998
Object recognition 1/2 sec 30 - 50 sziklai_studies_1956
Optimal performance in laboratory motor tasks $\sim$15 sec 10 - 12 fitts_information_1954, fitts_information_1964
Reading (English) min 28 - 45 rayner_eye_1998
Speech in 17 languages < 1 min 39 coupe_different_2019
Speed card 12.74 sec 17.7 internationalassociationofmemory_speed_2024
StarCraft (e-athlete) min 10  guinnessworldrecordslimited_most_2023
Tetris (Rank S) min $\sim 7$  tetrio_tetra_2024
Typing (English) min - hr $10$  dhakal_observations_2018, shannon_prediction_1951
Table 1: The information rate of human behaviors
How should one interpret a behavioral throughput of 10 bits/s? That number is ridiculously small compared to any information rate we encounter in daily life. For example, we get anxious when the speed of the home WiFi network drops below 100 megabits/s, because that might compromise our enjoyment of Netflix shows. Meanwhile, even if we stay awake during the show, our brain will never extract more than 10 bits/s of that giant bitstream. More relevant to the present arguments, the speed of human behavior is equally dwarfed by the capacity of neural hardware in our brains, as elaborated in the following section.
@@ 4 The information capacity of the nervous system
From the moment that Claude Shannon founded information theory shannon_mathematical_1948, neurophysiologists have applied that framework to signaling in the nervous system. For example, one can view a single neuron as a communication channel. In sensory systems, in particular, we are quite sure what the neurons are communicating about, namely the sensory stimulus, which also happens to be under the experimenter’s control. To measure how much information the neuron conveys, one proceeds as described above for behavior: (1) Identify the range of outputs of the neuron across all possible stimuli. (2) Determine what part of that range is “noise” vs “signal”; typically this involves repeating the same stimulus many times and measuring the response variation. (3) Use the expressions for entropy and mutual information to derive an information rate (see Appendix B).
For example, the photoreceptor cells in the human eye convert dynamic light input to a continuously varying membrane voltage. Using the approach described in Appendix B, one finds that one human cone photoreceptor can transmit information at $\approx 270$ bits/s. The population of 6 million cones in one eye has a capacity of about 1.6 gigabits/s. From that enormous bitstream, the brain of the typist sifts out just the 10 bits/s needed to perform the behavioral task correctly. To represent this degree of filtering, we define the dimension-less “sifting number”:
$Si=\textrm{Sifting Number}=\frac{\textrm{Sensory information rate}}{\textrm{Behavioral throughput}}\approx\frac{\textrm{1 Gbit/s}}{\textrm{10 bit/s}}=10^{8}$ (2)
Unlike photoreceptors, most neurons in the central nervous system communicate via action potentials. The message from one neuron to the next depends entirely on the timing of their spikes. Again, one can adapt the methods of information theory and ask how much information gets transmitted that way. The answer is about 2 bits per spike borst_information_1999. This holds across neurons from many different species and over a wide range of firing rates.
This ballpark number of 2 bits/spike allows us to estimate the information rate at different stages of sensory processing. For example, the output of the eye consists of 1 million axons from retinal ganglion cells kandel_low-level_2021. With particularly strong stimuli one can drive them at an average rate of 50 Hz. So the capacity of the optic nerve is about 100 megabits/s or less. Note this is 10 times smaller than the capacity of cone photoreceptors: The circuits inside the retina already compress the visual signal substantially, by at least a factor of 10.
The example of retinal ganglion cells suggests that when driven to high firing rates, individual neurons may transmit hundreds of bits per second. In the mammalian cortex, the average firing rate of a neuron is very low, typically a few spikes per second lennie_cost_2003. However, even under those conditions, a single neuron will transmit $\sim$10 bits/s – equivalent to the information throughput of the entire human being.
@@ 5 The paradox of slow behavior
To reiterate: Human behaviors, including motor function, perception, and cognition, operate at a speed limit of 10 bit/s. At the same time, single neurons can transmit information at that same rate or faster. Furthermore, some portions of our brain, such as the peripheral sensory regions, clearly process information dramatically faster.
Before approaching the paradox, we should agree on its magnitude, which sets an expectation for the magnitude of the answer. We need to explain a factor of $Si=10^{8}$ , namely the ratio between peripheral information processing and the overall throughput of the brain. Numerically speaking, this may be the largest unexplained number in brain science. This order of magnitude sets a filter for hypothetical solutions. Suppose, for example, that a particular hypothesis explains why people can respond two times faster in one task condition than in another. That might be enough to publish in a prominent journal, but the tiny effect size doesn’t even make a dent in the paradox.
@@ 5.1 What about photographic memory?
One scientific report stands out here, about a single human subject who could fuse a random dot stereogram after each monocular image had been presented on different days stromeyer_detailed_1970. Each image had $100\times 100$ pixels and was presented for 2 minutes, which would require a perceptual rate of 100 bits/s, ten times what we propose. Now a ten-fold discrepancy is nothing dramatic, given the million-fold conundrum we pose. Nevertheless, there are good reasons to question this claim. For example, neural signals from the two eyes are merged already in the primary visual cortex. So the subject’s brain would have to somehow inject the memorized image into neurons very early in the visual system, or perhaps even store it there. The claim is so unexpected and potentially revolutionary that it should have triggered many follow-up experiments. No supporting reports have appeared, suggesting that those studies did not go as hoped for.
@@ 5.2 But what about the rich detail everywhere in my visual scene?
However, this is an illusion, called “subjective inflation” in the technical jargon oregan_solving_1992, knotts_subjective_2019. People feel that the visual scene is sharp and colorful even far in the periphery because in normal life we can just point our eyes there and see vivid structure. In reality, a few degrees away from the center of gaze our resolution for spatial and color detail drops off drastically, owing in large part to neural circuits of the retina banks_peripheral_1991. You can confirm this while reading this paper: Fix your eye on one letter and ask how many letters on each side you can still recognize rayner_eye_1998. Another popular test is to have the guests at a dinner party close their eyes, and then ask them to recount the scene they just experienced. These tests indicate that beyond our focused attention, our capacity to perceive and retain visual information is severely limited, to the extent of “inattentional blindness” simons_gorillas_1999, cohen_what_2016.
@@ 5.3 But what about unconscious processing?
Another objection to a very low rate of perception is that, by definition, it only takes into account things that we can consciously experience and access, so that we can talk about them graham_what_1869. Surely, this argument goes, the myriad images that flit over our retina every day must carry a great deal of fine-grained information into the brain. Perhaps this information cannot be accessed consciously as photographic memory, but it is still stored in the brain.
@@
A specific proposal for unconscious processing involves the development of visual circuits. It is claimed that the visual system learns about the statistics of natural scenes from being exposed to millions of those images over time. Through mechanisms of synaptic plasticity bi_synaptic_2001, the structure of neural circuits responds to visual experiences in the course of early development, which alters the subsequent function of those circuits katz_synaptic_1996. How much information does this process extract from the environment?
@@
The famous stripe-rearing experiments on kittens can serve as an example stryker_physiological_1978. The kittens were raised in an environment painted with vertical black and white stripes. From the kitten’s perspective, because it still got to shake its head, the orientation of the stripes was restricted to perhaps $\pm$20 degrees from the vertical. This did lead to some reorganization of brain wiring in the visual cortex. Let us suppose that the kitten brain has completely incorporated the change in visual statistics. Whereas in the natural world, all orientations are about equally represented across 180 degrees, in the new environment they are restricted to a 40-degree range. The information associated with such a change in probability distributions is $\log_{2}(180/40)\approx 2$ bits.
@@
By perfectly learning the new distribution of angles, the kitten acquired only two bits. The same argument applies to other aspects of natural image statistics: The distribution of spatial and temporal frequencies differs from the uniform, but the shape of that distribution can be captured by just a few bits olshausen_natural_1996. This is also true of the distribution of color spectra. In each of these cases, processes of activity-dependent development can learn only a few bits over the animal’s lifetime.
@@
Why is the information about scene statistics so limited? The main reason is that they are translation-invariant: To good approximation the frequency spectrum of a natural image is the same in the bottom left as in the top right. Consequently, it is not necessary to learn separate statistics at each point in the visual field or to specify neural circuitry differently at each point. Nature “knows” about this translation-invariance and has exploited it in a clever way: by inventing families of neuronal cell types that are cloned across the visual field roska_retina_2014, sanes_types_2015. For example, the human retina contains many millions of neurons, but they come in only $\sim 100$ types grunert_cell_2020. Cells of a given type are distributed evenly across the retina. And so are the synaptic circuits between cells, because synapses are specified by type identity. In this way, modifying a single gene can alter synaptic circuitry in the same way at thousands of different sites within the visual field sanes_synaptic_2020. Modern image recognition systems have copied this invention in the form of convolutional neural networks, where all units within a layer share the same convolution kernel, thus requiring only a few parameters to be learned lecun_gradient-based_1998.
@@
Our nervous system uses this convolutional trick even when adapting in real-time to a change in image statistics. For example, electrical synapses in the retina can be modified by the neuromodulator dopamine bloomfield_diverse_2009. Thus a change in a single variable, namely the diffuse concentration of dopamine, modifies visual processing everywhere in the visual field. These adaptations occur on the scale of minutes, so again, only a small rate of information flow from the environment is needed for these updates.